# BigGait: Learning Gait Representation You Want by Large Vision Models

### 29 Feb 2024 v1

@@ Abstract
Gait recognition stands as one of the most pivotal remote identification technologies and progressively expands across research and industrial communities. However, existing gait recognition methods heavily rely on task-specific upstream driven by supervised learning to provide explicit gait representations, which inevitably introduce expensive annotation costs and potentially cause cumulative errors. Escaping from this trend, this work explores effective gait representations based on the all-purpose knowledge produced by task-agnostic Large Vision Models (LVMs) and proposes a simple yet efficient gait framework, termed BigGait. Specifically, the Gait Representation Extractor (GRE) in BigGait effectively transforms all-purpose knowledge into implicit gait features in an unsupervised manner, drawing from design principles of established gait representation construction approaches. Experimental results on CCPG, CAISA-B* and SUSTech1K indicate that BigGait significantly outperforms the previous methods in both self-domain and cross-domain tasks in most cases, and provides a more practical paradigm for learning the next-generation gait representation. Eventually, we delve into prospective challenges and promising directions in LVMs-based gait recognition, aiming to inspire future work in this emerging topic. The source code will be available at https://github.com/ShiqiYu/OpenGait.
@@ 1 Introduction
Gait recognition methods [3, 29, 51, 28, 27, 49] currently in use heavily depend on upstream tasks driven by supervised learning, as illustrated by Fig. 1. These tasks encompass a range of objectives, including, but not limited to, pedestrian segmentation [44], human parsing [31], body posture [2], and SMPL [32] model estimation. In general, the upstream model serves the purpose of providing task-specific priors to filter out the gait-unrelated elements within walking videos, especially for RGB-encoded background and texture characteristics. Next, some inductive biases may be utilized to refine the intermediate gait representation, ranging from basic operations like size alignment [41], coordinate normalization [15] to sophisticated ones like making the silhouette edges differentiable [28] and enforcing the appearance reconstruction [49, 27]. Thus existing gait representations are largely predetermined by supervised third-party tasks.
Escaping from the reliance on upstream task-specific outcomes, this paper makes a pioneering effort to build the desired gait representation based on task-agnostic knowledge acquired from Large Vision Models (LVMs). We are strongly motivated by the following insights:
@@
  • Self-supervised pre-training of LVMs obviates the need for labeled datasets required for training the state-of-the-art upstream supervised algorithms, thus alleviating the substantial costs necessary for annotating elements such as the silhouette, skeleton, and more, on a large scale.
  • The task-agnostic knowledge embedded in LVMs is learned from web-scale data collections without task-specific supervision, thus ideally avoiding cumulative errors imposed by specific upstream tasks to a large extent.
Figure 2: The illustration of a) a body proportion preservation method is provided in the Supplementary Material, and the visualization of the intermediate representation outed by b) mask, c) appearance, and d) denoising branch.
@@
Overall, the proposed GRE module learns the design principles of many established gait representations and distinguishes itself by not necessitating task-specific supervision with leveraging all-purpose knowledge. Technically, GRE relies solely on soft constraints to extract effective gait representations, i.e., capturing truthful body structural characteristics while effectively excluding gait-unrelated noise. Moreover, this work also finds that the utilization of soft constraints may introduce two major challenges: (1) Interpretability. In contrast to the more visually intuitive modality such as the silhouette and skeleton, which directly convey physical meanings, the learned representation derived from soft geometrical constraints, presented as the multi-channel feature map, may not offer the same level of intuitiveness. (2) Purity. Completely suppressing the texture characteristics of RGB videos through soft constraints remains a challenge for gait recognition [28]. Our comprehensive cross-clothing and cross-domain evaluations provide strong evidence of the dominant role of gait features within the acquired gait representations. Nevertheless, we find that data distribution may still influence outcomes.
@@
  • We establish the effectiveness of all-purpose knowledge for gait description, as evidenced by its discriminability and generalization across various gait datasets, including CCPG [25], CASIA-B [46] and SUSTech1K [38]. BigGait is the state of the art in most cases.
@@ 2 Related Works
Upstream Models for Gait Representation Learning. Walking videos often contain numerous extraneous elements, including background distractions and variations in color, texture, clothing, and carried items within the foreground. Prior methods have introduced a range of task-specific models to address these challenges. These models are designed to provide binary silhouettes [3, 10, 30, 13], body skeletons [29, 42], human parsing [51], SMPL models [27], or a combination of these representations [50]. These representations serve as the input for downstream gait models. Moreover, several end-to-end methods [49, 28] aim to optimize the intermediate gait representation globally. However, it is important to recognize that these upstream tasks heavily rely on supervised learning, giving rise to three primary challenges: a) considerable annotation costs, b) restricted form of gait representation, and c) potential accumulation of estimation errors. In this paper, BigGait instantiates the upstream model as a self-supervised LVM, constructing the desired gait representation based on the all-purpose knowledge in a novel manner.
@@
Large Vision Models. Inspired by the success of LLMs [7, 1], researchers have ventured into exploring foundation models in the field of computer vision [34, 23, 36, 36], i.e., learning generalizable features from web-scale data sources by scalable vision models. For example, CLIP [36] used a form of textual supervision to guide the training of visual representation. SAM [23] developed a promotable model and pre-trained it on a broad dataset using a task that enables powerful generalization for image segmentation. This paper focuses on another alternative termed self-supervised LVMs [5, 6, 19], where features are learned from images alone without supervision. Specifically, the employed DINOv2 [34] demonstrated that self-supervised learning has the potential to learn all-purposed visual features if pre-trained on a large quantity of curated data. Here, all-purpose features aim to work out of the box on any task, e.g., image classification, and dense recognition tasks like semantic segmentation and depth estimation. The main goal of this paper is to learn gait representation from these all-purpose features, thereby benefiting from the advantages of LVMs.
@@ 3.1 Overview
BigGait takes an RGB video as input and processes each frame in parallel. For simplicity, we focus on a single frame in the following description. Before feeding the image into the upstream model, this paper adopts a Pad-and-Resize method to resizing it into a fixed resolution of $448\times 224$, while keeping the original body proportions. The Pad-and-Resize method is provided in the Supplementary Material.
Upstream Model. The upstream DINOv2 owns a scalable ViT [8] as the backbone, with parameter counts ranging from 21M to 1,100M. In this study, we opt for the smallest and second largest counterparts, namely, ViT-S/14 (21M) and ViT-L/14 (300M), to construct BigGait-S and BigGait-L, respectively. Given the resized RGB image, the ViT first divides it into non-overlapping patches with a size of $14\times 14$, thus taking $32\times 16$ tokenized vectors as input. Next, the positional encodings are added, and subsequent stacked ViT blocks transform the features. We gather tokens from low to high layers with uniform intervals, and all these $32\times 16$ tokens will be concatenated according to the spatial correspondence to form a feature map. These details make no difference from the official implementation [34] thus Fig. 3 ignores them for brevity.
We leverage the officially provided model checkpoint pre-trained on LVD-142M [34] and freeze the imported parameters during gait metric learning. To reduce carbon emissions, we establish BigGait-S as the reference version, even though BigGait-L exhibits slightly superior performance. Thanks to this selection, the upstream part of BigGait is even smaller than many typical segmentation, human parsing, and pose estimation networks widely used by existing gait methods, as shown on Tab. 1.
Table 1: Parameter and GFLOPs of different upstream models. The GFLOPs are calculated for the input size $448\times 224$.
Type Methods Backbone #Params GFLOPs
Segmentation UNet3+ [20] UNet3+ 27.0M 308.2
DeepLabV3+ [4] ResNet-50 26.8M 43.7
Human Parsing SCHP [24] A-CE2P 66.6M 33.5
HRNet [44] HRNet-W48 70.1M 61.9
Pose Estimation HRNet [40] HRNet-W32 28.5M 15.8
DEKR [16] HRNet-W32 29.5M 17.1
All-purpose (Ours) DINOv2 [34] ViT-S/14 21.5M 11.0
ViT-L/14 302.9M 155.3
@@
Downstream Model. We make a minor modification to GaitBase [13], so that it can process the two-stream input. As showin in Fig. 3, $B_{1},B_{2},B_{3}$ and $B_{4}$ refer to the standard blocks within GaitBase, where $B_{1}^{\text{ap}}$ and $B_{1}^{\text{de}}$ own the different parameters but share the identical structure with $B_{1}$. Then, an attention block possessing the widely used cross-and-select architecture [26] targets aggregating the generated two-branch features. Excluding the above modifications, our implementation of GaitBase makes no architectural difference from the official one [13]. The training is driven by the triplet $L_{tri}$ and cross-entropy $L_{L_{ce}}$ losses.
Visualization. To understand the content of each intermediate feature map, namely $f_{c},f_{m},f_{ap}$, and $f_{de}$ depicted in Fig. 3, we conduct a PCA using the pixel-wise features from the entire dataset. Projecting pixel-wise features onto the first $3$ PCA bases can allow us to visualize a high-dim feature map as an RGB color image. This visualization approach is inherited from DINOv2 [34].
@@ 3.2 Gait Representation Extractor
As shown in Fig. 3, $f_{1},f_{2},f_{3},f_{4}$ denote the feature map generated by various stages of the ViT backbone with the corresponding semantic hierarchy spanning from low to high levels. Each of them is 2×2$\times$2 × upsampled by bilinear interpolation to improve the resolution, i.e., exhibited as a 3-D tensor with a size of $384\times 64\times 32$ while the first dimension denotes the output channel of the upstream DINOv2. Additionally, we concatenate these feature maps along the channel dimension to form $f_{c}$, i.e., with a size of $1,536\times 64\times 32$, which is rich in both low-level details and high-level semantics. As shown in Fig. 3, $f_{c}$ is dominated by the foreground and background regions accompanied by noisy spots.
Mask Branch. Inspired by the design of the gait silhouette, the GRE module develops a mask branch to remove the background noise. Specifically, an auto-encoder is proposed to generate the foreground mask based on $f_{4}$:
@@
Intuitively, the mask branch works like a PCA by searching two mutually exclusive first components. Thanks to the segmentable nature of $f_{4}$, the mask $m$ can be learned easily in this way. Notably, the $m$ is a two-channel tensor, and which channel represents the foreground is unknown. This paper selects the channel whose activations lay more in the image center as the foreground mask. Moreover, we perform the binarization and closing operations to reduce the potential cavities and breakpoints. These common tricks do not appear in Fig. 3 for brevity, i.e., the symbol of $m$ in Fig. 3 denotes the final foreground mask.
After masking the background regions, i.e., comparing the input $f_{c}$ with the output $f_{m}$, Fig. 3 shows that this process not only largely filters out the background noise but also makes foreground features diverse and discriminative over body parts, resulting in a parsing-like effect.
@@
By comparing the input $f_{m}$ with the output $f_{ap}$, as shown in Fig. 3, the emergence of noisy elements within the latter should be primarily driven by the downstream gait model rather than the linear $E_{\text{ap}}$. This observation is consistent with findings in related studies [28, 25], suggesting that the gait model tends to emphasize the clearly unchanged gait-unrelated cues rather than the subtle gait patterns. Further visualizations in Sec. 3.3 reveal the correlation between these noises with the clothing texture on the body.
@@
where $E_{\text{de}}$ denotes non-linear layers composed by a $1\times 1$ convolution, a batch normalization, a GELU, and an additional $1\times 1$ convolution. Its output channel is set to $C$ which is the same as that of the appearance branch for brevity. The softmax function is performed along the channel dimension to normalize the features. The smoothness loss is defined by the absolute value of channel-wise image gradients along both the width and height axes. In Eq. 3, $\text{sobel}_{x}$ and $\text{sobel}_{y}$ refer to the x-axis and y-axis Sobel operator [39], and the symbol of *** represents the convolution computation.
From a definitional standpoint, $L_{smo}$ can not remove the gradual textural noise effectively. But for gait recognition, we consider that the evident textures on body images, e.g., the clothing logo and pattern, are generally high-frequency characteristics that can be easily excluded by $L_{smo}$.
@@
Overall, the learning of the denoising branch is entirely driven by the above two soft constraints. These losses only perform in the foreground produced by the mask branch.
@@ 3.3 Visualization of Intermediate Representation
To provide an intuitive understanding, we visualize each intermediate representation in BigGait, and compare it with some traditional gait representations, as shown in Fig. 4. All-purpose features $f_{m}$ created by the upstream model (DINOv2) provide diverse and discriminative body parts after masking the background regions, i.e., purple head, red abdomen, green limbs and blue shoe. Both $f_{ap}$ in the appearance branch and $f_{de}$ in the denoising branch are derived from $f_{m}$, with the former inheriting features by a linear transformation and showing high-frequency texture noises, and the latter embodying highly consistent skeleton-like structure representation since deploying smoothness and diversity constraints to denoise most texture elements on the body. However, compared with the traditional gait representations with direct physical meanings, the gait representation based on soft geometrical constraints may need more interpretability on its physical meaning in the future.
Figure 4: The visualization of intermediate representations generated by BigGait v.s. three typical gait representations.

### 22 Mar 2024 v2

@@ Abstract
Gait recognition stands as one of the most pivotal remote identification technologies and progressively expands across research and industry communities. However, existing gait recognition methods heavily rely on task-specific upstream driven by supervised learning to provide explicit gait representations like silhouette sequences, which inevitably introduce expensive annotation costs and potential error accumulation. Escaping from this trend, this work explores effective gait representations based on the all-purpose knowledge produced by task-agnostic Large Vision Models (LVMs) and proposes a simple yet efficient gait framework, termed BigGait. Specifically, the Gait Representation Extractor (GRE) within BigGait draws upon design principles from established gait representations, effectively transforming all-purpose knowledge into implicit gait representations without requiring third-party supervision signals. Experiments on CCPG, CAISA-B* and SUSTech1K indicate that BigGait significantly outperforms the previous methods in both within-domain and cross-domain tasks in most cases, and provides a more practical paradigm for learning the next-generation gait representation. Finally, we delve into prospective challenges and promising directions in LVMs-based gait recognition, aiming to inspire future work in this emerging topic. The source code is available at https://github.com/ShiqiYu/OpenGait.
@@ 1 Introduction
Existing gait recognition methods [3, 29, 53, 28, 27, 50, 51, 31] heavily depend on upstream tasks driven by supervised learning, as illustrated by Fig. 1. These tasks encompass a range of objectives, including, but not limited to, pedestrian segmentation [45], human parsing [32], body posture [2], and SMPL [33] model estimation. In general, the upstream model serves the purpose of providing task-specific priors to filter out the gait-irrelevant cues within walking videos, especially for RGB-encoded background and texture characteristics. Next, some inductive biases may be utilized to refine the intermediate gait representation, ranging from basic operations like size alignment [42], coordinate normalization [15] to sophisticated ones like making the silhouette edges differentiable [28] and enforcing the appearance reconstruction [50, 27]. Briefly, existing gait representations are largely predetermined by upstream supervised tasks.
Escaping from the reliance on upstream task-specific models, this paper makes a pioneering effort to acquire desired gait features from task-agnostic Large Vision Models (LVMs). We are strongly driven by the following insights:
@@
  • Self-supervised pre-training of LVMs obviates the need for labeled datasets required for training the state-of-the-art gait upstream models, thus eliminating the substantial costs necessary for annotating elements such as the silhouette, skeleton, and more, on a large scale.
  • The task-agnostic knowledge embedded in LVMs is learned from web-scale datasets without task-specific supervision, thus ideally avoiding error accumulation imposed by specific upstream tasks to a large extent.
Figure 2: The illustration of (a) body proportion preservation trick is provided in the Supplementary Material. (b)-(d) respectively present the visualization of intermediate representation generated by (b) mask, (c) appearance, and (d) denoising branch.
@@
Overall, our GRE module adopts the design principles of some established gait representations and distinguishes itself by leveraging all-purpose knowledge and removing upstream task-specific supervision. Technically, GRE relies solely on soft constraints to extract effective gait representations, i.e., capturing truthful body structural characteristics while effectively excluding gait-irrelevant noise. Moreover, this work also finds that the utilization of soft constraints may introduce two major challenges: (1) Interpretability. In contrast to the more visually intuitive modality such as the silhouette and skeleton, which directly convey physical meanings, the learned representation derived from soft geometrical constraints, presented as the multi-channel feature map, may not offer the same level of intuitiveness. (2) Purity. Completely suppressing the texture characteristics of RGB videos through soft constraints remains a challenge for gait recognition [28]. Our comprehensive cross-clothing and cross-domain evaluations provide strong evidence of the dominant role of gait features within the acquired gait representations. Nevertheless, we find that data distribution may still influence outcomes.
@@
  • We establish the effectiveness of all-purpose knowledge for gait description, as evidenced by its discriminability and generalization across various gait datasets, including CCPG [25], CASIA-B* [47] and SUSTech1K [39]. BigGait achieves the SoTA performance in most cases.
@@ 2 Related Works
Upstream Models for Gait Representation Learning. Walking videos often contain numerous extraneous elements, including background distractions and variations in color, texture, clothing, and carried items within the foreground. Prior methods have introduced a range of task-specific models to address these challenges. These models are designed to provide binary silhouettes [3, 10, 30, 13], body skeletons [29, 43], human parsing [53], SMPL models [27], or a combination of these representations [52], to serve as the input for downstream gait models. Moreover, several end-to-end methods [50, 28] aim to optimize the intermediate gait representation globally. However, it is important to recognize that these upstream tasks heavily rely on supervised learning, giving rise to three primary challenges: a) considerable annotation costs, b) restricted form of gait representation, and c) potential accumulation of estimation errors. In this paper, BigGait instantiates the upstream model as a self-supervised LVM, constructing the desired gait representation by leveraging the all-purpose knowledge in a novel manner.
@@
Large Vision Models. Inspired by the success of LLMs [7, 1], researchers have ventured into exploring foundation models in the field of computer vision [35, 23, 37], i.e., learning generalizable features from web-scale data sources by scalable vision models. For example, CLIP [37] used a form of textual supervision to guide the training of visual representation. SAM [23] developed a promotable model and pre-trained it on a broad dataset using a task that enables powerful generalization for image segmentation. This paper focuses on another alternative termed self-supervised LVMs [5, 6, 19], where features are learned from images alone without supervision. Specifically, the employed DINOv2 [35] demonstrated that self-supervised learning has the potential to learn all-purposed visual features if pre-trained on a large quantity of curated data. Here, all-purpose features aim to work out of the box on any task, e.g., image classification, and dense recognition like semantic segmentation and depth estimation. The main goal of this paper is to learn gait representation from these all-purpose features, thereby benefiting from the advantages of LVMs.
@@ 3.1 Overview
BigGait takes an RGB video as input and processes each frame in parallel. For simplicity, we focus on a single frame in the following description. Before feeding the image into the upstream model, this work adopts a Pad-and-Resize trick to resize it into a fixed resolution of $448\times 224$ while keeping the truthful body proportions. The illustration of this detail is provided in the Supplementary Material.
Upstream Model. The upstream DINOv2 owns a scalable ViT [8] as the backbone, with parameter counts ranging from $21$M to $1,100$M. In this study, we opt for the smallest and second largest counterparts, namely, ViT-S/14 ($21$M) and ViT-L/14 ($302$M), to construct BigGait-S and BigGait-L, respectively. Given the resized RGB image, the ViT first divides it into non-overlapping patches with a size of $14\times 14$, thus taking $32\times 16$ tokenized vectors as input. Next, the positional encodings are added, and subsequent stacked ViT blocks transform the features. We gather tokens from low to high layers with uniform intervals, and all these $32\times 16$ tokens are concatenated according to the spatial correspondence to form a feature map. These details follow the official implementation [35]. Finally, we upsample the feature map to $64\times 32$ as output.
We leverage the officially provided model checkpoint pre-trained on LVD-142M [35] and freeze the imported parameters during gait metric learning. To reduce carbon emissions, we establish BigGait-S as the reference version, even though BigGait-L exhibits slightly higher performance. Thanks to this choice, the upstream part of BigGait is even smaller than many typical segmentation, human parsing, and pose estimation networks widely used by existing gait methods, as shown in Tab. 1.
Table 1: Parameter and GFLOPs of different upstream/overall methods. The GFLOPs are calculated for the input size $448\times 224$.
Upstream Obj. Methods Backbone #Params (M) GFLOPs
Segmentation UNet3+ [20] UNet3+ $27.0$ $308.2$
DeepLabV3+ [4] ResNet-50 $26.8$ $43.7$
Human Parsing SCHP [24] A-CE2P $66.6$ $33.5$
HRNet [45] HRNet-W48 $70.1$ $61.9$
Pose Estimation HRNet [41] HRNet-W32 $28.5$ $15.8$
DEKR [16] HRNet-W32 $29.5$ $17.1$
All-purpose (Ours) DINOv2 [35] ViT-S/14 $21.5$ $11.0$
ViT-L/14 $302.9$ $155.3$
Overall Upstream Downstream #Params (M) GFLOPs
Sil.-based DeepLabV3+ GaitBase $34.2$ $45.4$
Parsing-based SCHP GaitBase $74.0$ $35.2$
Skeleton-based HRNet-W32 Gait-TR $29.0$ $31.2$
BigGait (Ours) ViT-S/14 + GRE GaitBase $30.8$ $12.7$
@@
Downstream Model. We make a minor modification to GaitBase [13] so that it can process the two-stream input. As shown in Fig. 3, $B_{1},B_{2},B_{3}$ and $B_{4}$ refer to the standard blocks within GaitBase, where $B_{1}^{\text{ap}}$ and $B_{1}^{\text{de}}$ own the different parameters but share the identical structure with $B_{1}$. Then, an attention block possessing the widely used cross-and-select architecture [26] targets aggregating the generated two-branch features. Excluding the above modifications, our implementation of GaitBase makes no architectural difference from the official one [13]. The training is driven by the triplet $L_{tri}$ and cross-entropy $L_{L_{ce}}$ losses.
Visualization. To understand the content of each intermediate feature map, namely $f_{c},f_{m},f_{ap}$, and $f_{de}$ depicted in Fig. 3, we conduct a PCA using the pixel-wise features from the entire dataset. Projecting pixel-wise features onto the first three PCA bases allow us to visualize a high-dim feature map as an RGB color image. This visualization approach is inherited from DINOv2 [35].
@@ 3.2 Gait Representation Extractor
As shown in Fig. 3, $f_{1},f_{2},f_{3},f_{4}$ denote the feature map generated by various stages of the ViT backbone with the corresponding semantic hierarchy spanning from low to high levels, which is a common practice suggested by DINOv2. Each of them is 2×2$\times$2 × upsampled by bilinear interpolation to improve the resolution, i.e., exhibited as a 3-D tensor with a size of $384\times 64\times 32$ while the first dimension denotes the output channel of the upstream DINOv2. Additionally, we concatenate these feature maps along the channel dimension to form $f_{c}$, i.e., with a size of $1,536\times 64\times 32$, which is rich in both low-level details and high-level semantics. As shown in Fig. 3, $f_{c}$ is dominated by the foreground and background regions accompanied by noisy spots.
Mask Branch. Inspired by the design of the silhouette, the GRE module develops a mask branch to remove the background noise on the whole. Specifically, an auto-encoder is proposed to generate the foreground mask based on $f_{4}$:
@@
Intuitively, the mask branch works like a PCA by searching two mutually exclusive first components. Thanks to the segmentable nature of $f_{4}$, as shown in Fig. 4, the mask $m$ can be easily learned in this way. Notably, the $m$ is a two-channel tensor, and which channel represents the foreground is uncertain. So we select the channel whose activations lay more in the image center as the foreground mask. Moreover, we perform the binarization and closing operations to reduce the potential cavities and breakpoints. These common tricks do not appear in Fig. 3 for brevity, i.e., the symbol of $m$ in Fig. 3 denotes the final foreground mask.
After masking the background regions in $f_{c}$ with $m$, we obtain $f_{m}$:
$f_{m}=m\cdot f_{c},$ (2)
where $\cdot$ denote the multiplication. Fig. 3 shows that this process not only largely filters out the background noise but also makes foreground features diverse and discriminative over body parts, resulting in a parsing-like effect.
Figure 4: Visualization of unsupervised mask learning.
@@
By comparing the input $f_{m}$ with the output $f_{ap}$, as shown in Fig. 3, the emergence of noisy elements within the latter should be primarily driven by the downstream gait model rather than the linear $E_{\text{ap}}$. This observation is consistent with findings in related studies [28, 25], suggesting that the gait model tends to emphasize the clearly unchanged gait-irrelevant cues rather than the subtle gait patterns. Further visualizations in Sec. 3.3 reveal the correlation between these noises and the clothing texture on the body.
@@
where $E_{\text{de}}$ denotes non-linear layers composed by a $1\times 1$ convolution, a batch normalization, a GELU, and an additional $1\times 1$ convolution. Its output channel is set to $C$ which is the same as that of the appearance branch for brevity. The softmax function is performed along the channel dimension to normalize the features. The smoothness loss is defined by the absolute value of channel-wise image gradients along both the width and height axes. In Eq. 4, $\text{sobel}_{x}$ and $\text{sobel}_{y}$ refer to the x-axis and y-axis Sobel operator [40], and the symbol of *** represents the convolution operation.
From a definitional standpoint, $L_{smo}$ can not remove the gradual textural noise effectively. But for gait recognition, we consider that the evident textures on body images, e.g., the clothing logo and pattern, are generally high-frequency characteristics that can be easily eliminated by $L_{smo}$.
@@
Overall, the learning of the denoising branch is entirely driven by the above two soft constraints. These losses only perform in the foreground produced by the mask branch. Finally, we fuse $f_{ap}$ and $f_{de}$ using attention weights:
$f_{fusion}=Attn(\text{B}_{1}^{ap}(f_{ap}),\text{B}_{1}^{de}(f_{de})),$ (6)
where $Attn$ is an attention block, following [14].
@@ 3.3 Visualization of Intermediate Representation
To provide an intuitive understanding, we visualize each intermediate representation in BigGait, and compare it with traditional gait representations, as shown in Fig. 5. All-purpose features $f_{m}$ created by the upstream model (DINOv2) provide diverse and discriminative body parts after masking the background regions, i.e., purple head, red abdomen, green limbs and blue shoe. Both $f_{ap}$ in the appearance branch and $f_{de}$ in the denoising branch are derived from $f_{m}$, with the former inheriting features by a linear transformation and showing body shape representation with high-frequency texture noises, and the latter embodying highly consistent skeleton-like representation since deploying smoothness and diversity constraints removes most texture noises. However, compared with traditional gait representations with direct physical meanings, the gait representation based on soft geometrical constraints may need more interpretability on its physical meaning in future.
Figure 5: The visualization of intermediate representations generated by BigGait v.s. three traditional gait representations. The red boxes indicate regions with strong texture patterns.